### v2

# Title of the Paper

Abstract
This paper studies how payoff heterogeneity affects the value of information in beauty contest games. I show that public information is detrimental to welfare if and only if agents’ Katz-Bonacich centralities exhibit specific forms of heterogeneity, stemming from the network of coordination motives. A key insight is that agents may value the commonality of information so differently that some are harmed by their neighbors knowing what others know. Leveraging this insight, I also show that when the commonality of information is endogenously determined through information sharing, the equilibrium degree of information sharing can be inefficiently low, even without sharing costs.
Keywords: Beauty contest, coordination, value of information, network games.
1 Introduction
Many socioeconomic interactions are shaped simultaneously by incomplete information and complex strategic interdependencies. A user deciding on technology adoption often faces uncertainty about its intrinsic value and payoffs contingent on adoption patterns within their social network. An investor formulates investment strategies based on unknown economic fundamentals while anticipating the behavior of other market participants, especially influential ones. Within organizations, a worker facing task uncertainty often needs to coordinate their actions with specific individuals or teams, as dictated by formal and informal structures.
Despite the recognized prevalence of information and payoff heterogeneity, their interplay remains relatively underexplored. This paper bridges this gap by examining how payoff heterogeneity arising from networked interactions affects the welfare implications of public information. For this purpose, I introduce a beauty contest game incorporating a flexible form of coordination motives represented by a weighted and directed network. Agents incur quadratic losses from deviations of their actions from a common payoff state and the actions of their specific neighbors. Agents do not directly observe the state but receive private and public signals.
The main result of this paper provides a geometric characterization of interaction networks under which public information provision can be detrimental to welfare. This result might seem counterintuitive, as public information is typically beneficial because it informs an agent about the payoff state and what her neighbors know. However, public information also informs her neighbors about what the neighbors of her neighbors know. This increased commonality of information can distort her neighbors’ actions away from the state, thereby imposing a negative externality on her. Depending on the network topology, such negative externalities can dominate the informational gains from public information.
While the finding that public information can harm welfare in beauty contest games resonates with the seminal work of Morris and Shin, (2002), the underlying economic mechanism identified here is distinct. The anti-transparency result in Morris and Shin hinges on a specific welfare criterion independent of coordination losses; only deviations from the state affect aggregate welfare. The result in this paper requires no such assumption on the welfare function; instead, the detrimental effect of public information necessitates a sufficient degree of heterogeneity in coordination motives.
The formal analysis employs a linear-quadratic-Gaussian framework. Linear best responses, combined with Gaussian signals, ensure the existence of a unique linear equilibrium. In this equilibrium, the weight placed on the public signal is proportional to an agent’s Katz-Bonacich centrality, which captures her direct and indirect coordination motives (i.e., coordination motives of neighbors, neighbors’ neighbors, etc.). Thus, more central agents are more responsive to the public signal than less central ones in equilibrium.
The first result shows that public information provision decreases an individual agent’s equilibrium payoff if and only if the agent’s Katz-Bonacich centrality is sufficiently low relative to an aggregate measure of her neighbors’ centralities. For example, in core-periphery networks, harm from public information typically arises for peripheral agents who intend to coordinate with agents in a highly cohesive core. The core agents place substantial weight on public information due to their strong coordination motives, sometimes at the expense of aligning their actions with the payoff state. [^1] This observation is reminiscent of the anti-transparency result of Morris and Shin, (2002). This deviation by the core agents, in turn, makes it more difficult for the peripheral agents to simultaneously align their actions with the state and the core agents’ actions.
Public information may also decrease aggregate welfare. These detrimental welfare effects tend to arise in networks characterized by a few highly central agents serving as coordination hubs for many less-central agents. In core-periphery networks with a small, cohesive core and a large periphery, public information benefits core agents while potentially harming peripheral agents. Aggregate welfare declines if the losses incurred by the peripheral agents exceed the gains accruing to the core agents.
Finally, applying insights from the main analysis, I find a novel source of inefficiency in endogenous information sharing. Consider an extension where an agent initially possesses a signal with the potential for public disclosure and can commit to sharing it publicly. In this setting, the agent may choose not to share the signal in equilibrium even when its public disclosure would be socially desirable. In core-periphery networks, a peripheral agent may prefer to withhold her signal to prevent core agents from being overly responsive should the signal be shared. Such a decision by the peripheral agent can lead to substantial welfare losses, as the potential benefits to core agents are forgone.
2 Literature
This paper contributes to the literature on the social value of information. In their seminal work, Morris and Shin, (2002) developed the beauty contest model and showed that public information can harm welfare. Their model has been applied to diverse contexts, including technological investment (Angeletos and Pavan,, 2004), asset pricing (Allen et al.,, 2006), and political leadership (Dewan and Myatt,, 2008). Angeletos and Pavan, (2007) and Ui and Yoshizawa, (2015) characterized the social value of public information in more general linear-quadratic games. Other notable extensions include work by Colombo et al., (2014) on endogenous acquisition of private information, Bergemann and Morris, (2013) on optimal Bayes correlated equilibria, and Huo and Pedroni, (2020) on dynamic information. Pavan and Vives, (2015) and Angeletos and Lian, (2016) provide an excellent overview of this literature.
Almost all existing work in this literature assumes no payoff heterogeneity. [^2] There are a few exceptions. The solution method for dynamic beauty contest games developed by Huo and Pedroni, (2020) applies to network games. Recent work by Miyashita and Ui, (2023) proposed a semi-definite programming approach to information design problems in network games. Also, Morris and Shin’s anti-transparency result hinges on specific assumptions about the welfare function, particularly its independence from coordination losses (Angeletos and Pavan,, 2004; Hellwig,, 2005; Woodford,, 2005). In contrast, I study environments where agents have heterogeneous coordination motives and the welfare measure incorporates coordination losses. The detrimental effect of public information identified in this paper arises from a novel mechanism that necessitates payoff heterogeneity.
The paper also contributes to the literature on endogenous information acquisition within the beauty contest framework. Hellwig and Veldkamp, (2009) and Myatt and Wallace, (2012) developed flexible yet tractable information acquisition models where information publicity and precision are endogenously determined. More recently, Herskovic and Ramos, (2020) studied information acquisition from neighbors in endogenously formed networks.
A key feature of these models is that “agents want to know what others know” due to their coordination motives; specifically, the marginal benefit of observing a signal increases with the number of other observers. This is not necessarily true when coordination motives are heterogeneous, as agents may not want specific other agents to observe their signal. Using a simple model of endogenous information sharing, I show that such non-monotonicity can cause inefficiency in the equilibrium degree of information sharing.
A few recent papers have studied endogenous information acquisition in network games. Myatt and Wallace, (2019) introduced heterogeneous coordination motives to Myatt and Wallace, (2012). Leister, (2020) examined endogenous acquisition of private information in the network game of Ballester et al., (2006). [^3] In related models, Calvó-Armengol and de Martí, (2009), Calvó-Armengol et al., (2015), and Myatt and Wallace, (2018) also consider endogenous information acquisition with payoff heterogeneity. Their main focus is to characterize how equilibrium information acquisition depends on the network of payoff interdependencies, which is complementary to but different from mine. [^4] For instance, Proposition 2 of Myatt and Wallace, (2019) provides an equilibrium characterization similar to one in this paper: Agents with higher Katz-Bonacich centrality are more responsive than those with lower centrality to relatively public signals (whose publicity is endogenous in their model).
Finally, in a companion paper (Murayama,, 2025), I generalized the anti-transparency result of Morris and Shin, (2002) to large network games. That paper shows that when welfare solely depends on the action-state distance, public information provision harms welfare if and only if the mean square of the Katz-Bonacich centrality is sufficiently large. Interestingly, these detrimental effects may arise even when agents’ coordination motives are positive but arbitrarily small on average, highlighting the importance of indirect coordination motives.
3 Model
I introduce a beauty contest game in which agents’ coordination motives are represented as a weighted and directed network.
Payoff Function. There are $n$ agents, indexed by $i\in N\equiv\{1,\ldots,n\}$, where $n\geq 2$. Agent $i$’s action is $a_{i}\in\mathbb{R}$, and $a=(a_{i})_{i\in N}$ denotes an action profile. There is a common payoff state $\theta\in\mathbb{R}$. Each agent $i$’s payoff function is
$u_{i}(a,\theta)=-\left(1-\sum_{j\neq i}g_{ij}\right)(a_{i}-\theta)^{2}-\sum_{j\neq i}g_{ij}(a_{i}-a_{j})^{2}.$ (1)
The first term in (1) represents the loss from the deviation of agent $i$’s action from the payoff state. The second term is a sum of losses from deviations of $i$’s action from those of other agents. The coefficient $g_{ij}\geq 0$ measures the strength of $i$’s coordination motive with $j$. Throughout the paper, I assume that $\sum_{j\neq i}g_{ij}<1$ for all $i\in N$. An $n\times n$ matrix $G=[g_{ij}]$ collects these coordination motives, where $g_{ii}=0$ for all $i\in N$. This matrix $G$ can be viewed as the adjacency matrix of a weighted and directed network and is referred to as an interaction network. I call $d^{out}_{i}\equiv\sum_{j\neq i}g_{ij}$ $i$’s out-degree and call $j$ a neighbor of $i$ if $g_{ij}>0$.
Information Structure. The information structure is standard, following Morris and Shin, (2002). The payoff state $\theta$ follows an improper uniform prior distribution over the real line. Each agent $i$ observes a signal vector $\mathbf{s}_{i}=(x_{i},y)$, where $x_{i}$ is a private signal observed only by $i$, and $y$ is a public signal observed by all agents. Specifically, I define $x_{i}\equiv\theta+\varepsilon_{i}$ and $y\equiv\theta+\varepsilon_{0},$ where $\varepsilon_{i}$ and $\varepsilon_{0}$ are normally distributed with
$\mathbb{E}[\varepsilon_{i}]=\mathbb{E}[\varepsilon_{0}]=0,\mathrm{Var}[\varepsilon_{i}]=\tau_{x}^{-1},\mathrm{Var}[\varepsilon_{0}]=\tau_{y}^{-1},$
and the noise terms and $\theta$ are all mutually independent. The relative precision of private signals is denoted by $\gamma\equiv\tau_{x}/(\tau_{x}+\tau_{y})\in(0,1)$.
Equilibrium. The game defined above has a unique linear Bayesian Nash equilibrium. Since the payoff function (1) is strictly concave in $a_{i}$, the first-order condition is necessary and sufficient for equilibrium. Each agent $i$’s first-order condition is
$a_{i}=\left(1-d_{i}^{out}\right)\mathbb{E}_{i}[\theta]+\sum_{j\neq i}g_{ij}\mathbb{E}_{i}[a_{j}],$ (2)
where $\mathbb{E}_{i}[\cdot]\equiv\mathbb{E}[\cdot\mid\mathbf{s}_{i}]$ is the conditional expectation operator, conditional on $i$’s information $\mathbf{s}_{i}$. Consider linear strategies of the form $a_{i}=b_{i}y+(1-b_{i})x_{i}$ for all $i\in N$ [^5] Focusing on this form of linear strategies is without loss of generality. In Section 6.1, I show that the equilibrium slopes must sum to one for all agents in any linear equilibrium. Plugging these linear strategies into (2) yields
$b_{i}y+(1-b_{i})x_{i}=\left(1-\gamma+\gamma\sum_{j\neq i}g_{ij}b_{j}\right)y+\left(\gamma-\gamma\sum_{j\neq i}g_{ij}b_{j}\right)x_{i}.$ (3)
By matching the coefficients in (3), the equilibrium slopes must satisfy
$b_{i}-\gamma\sum_{j\neq i}g_{ij}b_{j}=1-\gamma.$ (4)
In matrix notation, these simultaneous equations are written as
$(I-\gamma G)\mathbf{b}=(1-\gamma)\mathbf{1},$ (5)
[…]
Proposition 1.
There exists a unique linear equilibrium in which agent $i$’s strategy is
$a^{*}_{i}=b^{*}_{i}y+(1-b^{*}_{i})x_{i},$ (6)
where $b^{*}_{i}=(1-\gamma)c_{i}(\gamma,G)$ is proportional to $i$’s Katz-Bonacich centrality.
In equilibrium, agent $i$’s direct and indirect coordination motives are positively associated with her responsiveness to the public signal. Intuitively, when $i$’s coordination motive strengthens, the commonality of the public signal becomes more valuable for aligning her action with her neighbors’ actions. Moreover, if $i$’s neighbors have stronger coordination motives, they will respond more to the public signal, further increasing the value of the public signal as a coordination device. Thus, not only an agent’s own coordination motive but also those of her neighbors (and neighbors’ neighbors, etc.) positively affect her responsiveness to the public signal. [^7] The relative precision of private information, $\gamma$, acts as the equilibrium discount factor in the Katz-Bonacich centrality. This role arises from how different orders of agents’ expectations about $\theta$ (e.g., $\mathbb{E}_{i}[\theta]$ and $\mathbb{E}_{i}[\mathbb{E}_{j}[\theta]])$ respond to changes in $\gamma$. A more detailed intuition is provided in Murayama, (2025).
4 Results
The main results show that public information provision can be detrimental to an individual agent and aggregate welfare, depending on the geometry of interaction networks. Most proofs and derivations are provided in Section 6.
4.1 Individual Payoff
[…]
$\mathbb{E}[X^{2}_{i}]$ $=\tau_{x}^{-1}(1-\gamma)c_{i}(\gamma,G)\left(c_{i}(\gamma,G)-2\right)+\tau_{x}^{-1};$ (7)
$\mathbb{E}[X_{i}X_{j}]$ $=\tau_{y}^{-1}(1-\gamma)^{2}c_{i}(\gamma,G)c_{j}(\gamma,G);$ (8)
$\sum_{j\neq i}g_{ij}\mathbb{E}[X_{i}X_{j}]$ $=\tau_{x}^{-1}(1-\gamma)c_{i}(\gamma,G)\left(c_{i}(\gamma,G)-1\right).$ (9)
The last equality (9) follows from the self-referentiality of the Katz-Bonacich centrality:
$c_{i}(\gamma,G)=1+\gamma\sum_{j\neq i}g_{ij}c_{j}(\gamma,G).$ (10)
Using these expressions, $i$’s equilibrium payoff simplifies to
$U_{i}(\tau_{x},\tau_{y},G)$ $=-\left(1-d_{i}^{out}\right)\mathbb{E}[X^{2}_{i}]-\sum_{j\neq i}g_{ij}\mathbb{E}[(X_{i}-X_{j})^{2}]$
$=-\mathbb{E}[X^{2}_{i}]-\sum_{j\neq i}g_{ij}\mathbb{E}[X_{j}^{2}]+2\sum_{j\neq i}g_{ij}\mathbb{E}[X_{i}X_{j}]$
$=\tau_{x}^{-1}(1-\gamma)\left(c_{i}^{2}(\gamma,G)-\sum_{j\neq i}g_{ij}c_{j}(\gamma,G)\left(c_{j}(\gamma,G)-2\right)\right)-\tau_{x}^{-1}(1+d_{i}^{out}).$ (11)
Let $\Delta U_{i}(\tau_{x},\tau_{y},G)\equiv U_{i}(\tau_{x},\tau_{y},G)-U_{i}(\tau_{x},0,G)$ be the effect of public information provision on agent $i$ [^9] Abusing notation, I write $U_{i}(\tau_{x},0,G)=-\tau_{x}^{-1}(1+d_{i}^{out})$ for agent $i$’s equilibrium payoff in the absence of public information. The following result is immediate from (11).
Proposition 2.
$\Delta U_{i}(\tau_{x},\tau_{y},G)<0$ if and only if
$c_{i}^{2}(\gamma,G)<\sum_{j\neq i}g_{ij}c_{j}(\gamma,G)(c_{j}(\gamma,G)-2).$ (12)
Proposition 2 states that the provision of public information is detrimental to agent $i$ if and only if, despite engaging in some coordination, her centrality is sufficiently low relative to an aggregate measure of her neighbors’ centralities. Note that inequality (12) never holds when the interaction network is regular (i.e., $d_{i}^{out}=d$ for all $i\in N$). [^10] In regular interaction networks, each agent’s centrality is given by $1/(1-\gamma d)$. Hence, the right-hand side of (12) must be smaller than the squared centrality. Thus, at least some heterogeneity in agents’ out-degrees is necessary for such a detrimental effect on individual payoff. The following example illustrates interaction networks that admit these adverse effects.
Example 1.
There are three agents, Ann, Bob, and Carol, in a team. Bob and Carol know each other well and intend to coordinate their action with each other. Ann is new to the team and intends to follow Bob’s action, while Bob and Carol do not intend to follow Ann’s action. Formally, the interaction network is given by $g_{AB}=\alpha$, $g_{BC}=g_{CB}=\beta$, and $g_{AC}=g_{BA}=g_{CA}=0$ (see Figure 1).
Ann Bob Carol $\alpha$$\beta$$\beta$
Figure 1: Interaction network among Ann, Bob, and Carol.
In this example, Ann may suffer from public information. First, Bob’s centrality is $c_{B}(\gamma,G)=1/(1-\gamma\beta)$, while Ann’s centrality is $c_{A}(\gamma,G)=(1+\gamma(\alpha-\beta))/(1-\gamma\beta)$. Plugging these into (12) and multiplying it by $(1-\gamma\beta)^{2}$ yields
$f(\alpha,\beta,\gamma)\equiv(1+\gamma(\alpha-\beta))^{2}-\alpha(2\gamma\beta-1)<0.$ (13)
Let $\mathcal{G}(\gamma)\equiv\{(\alpha,\beta)\in[0,1)^{2}:f(\alpha,\beta,\gamma)<0\}$ be the set of interaction parameters $(\alpha,\beta)$ under which Ann suffers from public information, given $\gamma$. It can be shown that $\mathcal{G}(\gamma)$ is nonempty if and only if $\gamma>(1+\sqrt{5})/4\approx 0.81$, and $\mathcal{G}(\gamma)\subseteq\mathcal{G}(\gamma^{\prime})$ for $\gamma\leq\gamma^{\prime}$ (see Figure 2). In Figure 2, Ann suffers from public information when $\beta$ is sufficiently large and $\alpha$ is relatively small but not too small.
(a) $\gamma=0.85$
(b) $\gamma=0.9$
(c) $\gamma=0.95$
Figure 2: $\mathcal{G}(\gamma)$ for $\gamma\in\{0.85,0.9,0.95\}$, where $x$-axis is $\alpha\in[0,1]$ and $y$-axis is $\beta\in[0.8,1]$.
The intuition for Proposition 2 is illustrated using Example 1. One might think that Proposition 2 is counterintuitive because Ann wants to align her action with $\theta$ and Bob’s action, and the public signal is informative about them. However, the public signal also informs Bob about what Carol knows (and vice versa), and Ann may suffer from Bob’s and Carol’s increased knowledge about what their respective neighbor knows. To see this, compare the current information structure $\mathcal{I}$ with an alternative information structure $\mathcal{I}^{\prime}$. Under $\mathcal{I}^{\prime}$, Ann and Bob observe the same signals as $\mathcal{I}$, while Carol observes another private signal $z\equiv\theta+\varepsilon_{z}$ instead of $y$. The noise term $\varepsilon_{z}$ follows a normal distribution with $\mathbb{E}[\varepsilon_{z}]=0$ and $\mathrm{Var}[\varepsilon_{z}]=\tau_{y}^{-1}$ and is independent of $\theta$ and all other noise terms. Thus, relative to $\mathcal{I}$, agents remain equally informed about $\theta$, Ann and Bob still observe the common signal $y$, but the commonality of information between Bob and Carol is diminished as Carol no longer observes $y$. The unique linear equilibrium under $\mathcal{I}^{\prime}$ is
$a^{\prime}_{A}$ $=(1-\gamma)(1+\gamma\alpha)y+\gamma(1-\alpha+\gamma\alpha)x_{A};$ (14)
$a^{\prime}_{B}$ $=(1-\gamma)y+\gamma x_{B};$ (15)
$a^{\prime}_{C}$ $=(1-\gamma)z+\gamma x_{C}.$ (16)
Comparing Ann’s equilibrium payoff under $\mathcal{I}$ and $\mathcal{I}^{\prime}$ yields that Ann is strictly better off under $\mathcal{I}^{\prime}$ if and only if
$f(\alpha,\beta,\gamma)<(1-\gamma\beta)^{2}\left((1+\gamma\alpha)^{2}+\alpha\right).$ (17)
Because the right-hand side of (17) is strictly positive, it follows that Ann suffers from public information only when she suffers from Bob and Carol knowing what their neighbor knows.
The detrimental effect on Ann, stemming from the increased commonality of a signal between Bob and Carol, is reminiscent of a mechanism highlighted by Morris and Shin, (2002). Morris and Shin show that when agents’ common coordination motive (analogous to $\beta$ in my model) is strong enough and the relative precision of private information is sufficiently high, public information provision increases the action-state distance, i.e., $\mathbb{E}[(a^{*}_{i}-\theta)^{2}]$. Similarly, the increased commonality of information between Bob and Carol can make Bob’s equilibrium action distant from $\theta$ [^11] Intuitively, this occurs because Bob and Carol’s coordination motive is so large that they choose to be much more responsive to the public signal than they would be when their objectives are to align their actions only with $\theta$. This is not desirable for Ann, whose objective is to simultaneously align her action with $\theta$ and Bob’s action.
In sum, public information becomes detrimental to Ann when the loss from Bob’s action being more distant from $\theta$ outweighs the benefit from having more precise information about $\theta$ and Bob’s action. The former loss is more prominent as Bob’s coordination motive with Carol increases. Moreover, Ann’s coordination motive with Bob should be relatively small but not too small for the detrimental effect; if it is too small (resp. too large), the benefit from knowing more about $\theta$ (resp. about Bob’s action) becomes dominant.
4.2 Aggregate Welfare
[…]
Proposition 3.
$\Delta W(\tau_{x},\tau_{y},G)<0$ if and only if
$\sum_{i\in N}c_{i}(\gamma,G)\left((1-d^{in}_{i})c_{i}(\gamma,G)+2d^{in}_{i}\right)<0,$ (18)
where $d^{in}_{i}\equiv\sum_{j\neq i}g_{ji}$ is agent $i$’s in-degree.
For inequality (18) to hold, at least one agent $i$ must have an in-degree $d^{in}_{i}>1$ and a Katz-Bonacich centrality $c_{i}(\gamma,G)>2d^{in}_{i}/(d^{in}_{i}-1)$. This condition requires stronger payoff heterogeneity, as neither regular nor undirected interaction networks satisfy it. [^13] If $G$ is undirected (i.e., $g_{ij}=g_{ji}$ for all $i,j\in N$), $d^{in}_{i}=d^{out}_{i}<1$ for all $i\in N$ by assumption. The following example illustrates interaction networks that exhibit such adverse welfare effects.
Example 2.
Consider a directed core-periphery network: $n$ agents are divided into $l\geq 2$ core agents and $m\geq 1$ peripheral agents, so $l+m=n$. Both core and peripheral agents intend to coordinate only with core agents; that is, $g_{ij}>0$ only if $j$ is a core agent. For simplicity, assume $d_{i}^{out}=\alpha$ if $i$ is peripheral and $d_{i}^{out}=\beta$ if $i$ is core. Example 1 is a special case of these core-periphery interaction networks with $l=2$ and $m=1$. Applying (18) to a core-periphery interaction network yields
$f(\alpha,\beta,\gamma)<-\frac{l}{m}(1-\beta(2\gamma\beta-1)).$ (19)
Let $\mathcal{H}(\gamma,l/m)\equiv\{(\alpha,\beta)\in[0,1)^{2}:f(\alpha,\beta,\gamma)<-l(1-\beta(2\gamma\beta-1))/m\}$ be the set of interaction parameters $(\alpha,\beta)$ that admit the negative welfare effects of public information, given $\gamma$ and $l/m$. It can be shown that $\mathcal{H}(\gamma,l/m)$ expands and converges to $\mathcal{G}(\gamma)$ as the number of peripheral agents relatively increases, i.e., $\mathcal{H}(\gamma,q)\subseteq\mathcal{H}(\gamma,p)$ for $p\leq q$ and $\mathcal{H}(\gamma,l/m)\rightarrow\mathcal{G}(\gamma)$ as $l/m\rightarrow 0$ (see Figure 3).
(a) $l/m=0.5$
(b) $l/m=0.2$
(c) $l/m=0.1$
Figure 3: $\mathcal{H}(\gamma,l/m)$ for $\gamma=0.9$ and $l/m\in\{0.5,0.2,0.1\}$.
The basic intuition for Proposition 3 parallels that for Proposition 2. One addition is that in Example 2, the in-degree of core agents reflects the extent to which peripheral agents target them for coordination. Distortion of core agents’ actions from the payoff state, resulting from their response to public information, imposes a larger aggregate loss when many peripheral agents attempt to coordinate with these distorted actions. Thus, for the total loss to peripheral agents to outweigh the total gain accrued by the core agents, at least some core agents must have a sufficiently large in-degree.
Finally, when the interaction network is regular, the welfare effect of public information increases with agents’ common out-degree. [^14] Suppose $d_{i}^{out}=d$ for all $i\in N$. Then, the term $c_{i}^{2}(\gamma,G)-\sum_{j\neq i}g_{ij}c_{j}(\gamma,G)\left(c_{j}(\gamma,G)-2\right)$ simplifies to $(1+d-2\gamma d^{2})/(1-\gamma d)^{2}$. Differentiating this ratio with respect to $d$ yields $((1-\gamma d)+2\gamma(1-d))/(1-\gamma d)^{3}$, which is strictly positive for all $d\in[0,1)$. Another implication of Proposition 3 (and Example 2) is that this monotonicity does not necessarily hold for more general interaction networks. The following proposition provides a class of information structures under which the welfare effect of public information can decrease as agents’ coordination motives become stronger.
Proposition 4.
[…]
4.3 Implications on Endogenous Information Sharing
The obtained results have implications for the efficiency of endogenous information sharing. Consider an extended model in which agent 1, initially possessing a signal $y$, chooses whether to make $y$ public or to keep it private before signal realizations. Agent $1$’s choice selects the information structure of the beauty contest game: If $y$ is made public, the information structure will be $\mathcal{I}$ as in the main analysis; if $y$ is kept private, the information structure will be $\mathcal{I}^{\dagger}$ under which agent 1 observes two private signals, $x_{1}$ and $y$, while each other agent $i\neq 1$ observes only her private signal $x_{i}$. [^16] It is implicitly assumed that other signals, such as $x_{1}$, are not shareable. One might justify this assumption by considering $x_{1}$ as tacit knowledge (e.g., personal experience) and $y$ as explicit knowledge (e.g., an analyst report). There is no cost to share information.
The two-stage game is solved by backward induction. The analysis requires comparing equilibrium payoffs under public sharing with those under private retention of the signal. As derived in Section 6, the equilibrium actions in the subgame following a choice to keep the signal private are
$a^{\dagger}_{1}=(1-\gamma)y+\gamma x_{1},\quad\text{and}\quad a^{\dagger}_{i}=x_{i}\ \text{for each }i\neq 1.$ (20)
Agent $1$’s equilibrium payoff and aggregate welfare are then computed as
$U^{\dagger}_{1}(\tau_{x},\tau_{y},G)$ $=\tau_{x}^{-1}(1-\gamma)+U_{1}(\tau_{x},0,G);$ (21)
$W^{\dagger}(\tau_{x},\tau_{y},G)$ $=\tau_{x}^{-1}(1-\gamma)(1+d_{1}^{in})+W(\tau_{x},0,G).$ (22)
Combining these expressions with Proposition 2 and 3 yields the following result.
Proposition 5.
The equilibrium degree of information sharing is inefficiently low, i.e., $U_{1}(\tau_{x},\tau_{y},G)<U^{\dagger}_{1}(\tau_{x},\tau_{y},G)$ and $W^{\dagger}(\tau_{x},\tau_{y},G)<W(\tau_{x},\tau_{y},G)$, if and only if the following two conditions hold:
$c_{1}^{2}(\gamma,G)-\sum_{j\neq 1}g_{1j}c_{j}(\gamma,G)\left(c_{j}(\gamma,G)-2\right)<1;$ (23)
$1+d^{in}_{1}<\sum_{i\in N}c_{i}(\gamma,G)\left((1-d_{i}^{in})c_{i}(\gamma,G)+2d_{i}^{in}\right).$ (24)
In words, the equilibrium degree of information sharing becomes inefficiently low if and only if: (i) agent 1 is harmed by making her information public, which typically occurs if her Katz-Bonacich centrality is sufficiently small relative to an aggregate measure of her neighbors’ centralities, and (ii) aggregate welfare would nonetheless increase if the information were shared, which typically occurs if the interaction network does not have agents with very high in-degrees and centralities. For instance, consider the core-periphery interaction networks introduced in Example 2. Assuming that agent 1 is peripheral, inequalities (23) and (24) simplify to
$f(\alpha,\beta,\gamma)<(1-\beta\gamma)^{2}<mf(\alpha,\beta,\gamma)+l(1-\beta(2\gamma\beta-1)).$ (25)
Let $\mathcal{J}(\gamma,l,m)\equiv\{(\alpha,\beta)\in[0,1)^{2}:f(\alpha,\beta,\gamma)<(1-\beta\gamma)^{2}<mf(\alpha,\beta,\gamma)+l(1-\beta(2\gamma\beta-1))\}$ be the set of interaction parameters $(\alpha,\beta)$ that admit inefficiently low equilibrium information sharing, given $\gamma$, $l$, and $m$. Because agent 1’s preference over information structures is “more aligned” with peripheral agents than with core agents, $\mathcal{J}(\gamma,l,m)$ shrinks as the number of peripheral agents relatively increases, i.e., $\mathcal{J}(\gamma,n-m^{\prime},m^{\prime})\subseteq\mathcal{J}(\gamma,n-m,m)$ for $m\leq m^{\prime}$ (see Figure 4). [^17] The formal argument is as follows. Fix $\gamma$ and take any $(\alpha,\beta)$ that satisfies $f(\alpha,\beta,\gamma)<(1-\beta\gamma)^{2}$. The right-hand side of (25) is linear in $m$ with slope $f(\alpha,\beta,\gamma)-(1-\beta(2\gamma\beta-1))$ by substituting $l=n-m$. This slope is bounded from above by $f(\alpha,\beta,\gamma)-1+2\gamma\beta^{2}-\beta$ $\leq f(\alpha,\beta,\gamma)-1+2\gamma\beta-\beta$ $\leq f(\alpha,\beta,\gamma)-1+2\gamma\beta-\gamma\beta$ $=f(\alpha,\beta,\gamma)-1+\gamma\beta$ $<0.$ The last strict inequality follows from the initial assumption. Thus, the right-hand side of (25) strictly decreases with $m$ whenever $f(\alpha,\beta,\gamma)<(1-\beta\gamma)^{2}$, which proves the claim.
(a) $m=44$
(b) $m=55$
(c) $m=60$
Figure 4: $\mathcal{J}(\gamma,l,m)$ for $\gamma=0.9$, $n=66$, and $m\in\{44,55,60\}$.
Intuitively, an agent may withhold her information to prevent her neighbors from excessively responding to her signal due to their strong coordination motives. Of course, keeping information private is not necessarily socially desirable, especially when other agents would benefit from the commonality of information. Thus, the resulting inefficiency can be substantial in the core-periphery interaction networks when the core is relatively large. This finding presents an interesting contrast to the case of public information provision, where the adverse welfare effects can be pronounced when the periphery is relatively large.
5 Conclusion
This paper studies how payoff heterogeneity affects the social value of information in beauty contest games where a network of coordination motives represents agents’ interactions. The analysis provides a geometric characterization of interaction networks under which public information provision can harm aggregate welfare. A key mechanism identified is that agents with relatively low Katz-Bonacich centrality may be harmed by public information because their highly central neighbors, by excessively responding to the public signal, distort their own actions relative to the payoff state. Building on this insight, the paper also highlights a novel channel through which information sharing becomes inefficient in beauty contest games: an agent may choose to withhold her signal to avoid the adverse consequences of her neighbors’ altered actions that would result from the increased commonality of information.
This paper is an initial step in understanding the social value of information with ex-ante heterogeneous agents. Promising avenues for future research include identifying optimal information dissemination policies, introducing endogenous information acquisition, and extending the analysis to more general linear-quadratic games.
6 Proofs
6.1 Proposition 1
The remaining part of the proof is to show that the sum of the equilibrium slopes must be one in any linear equilibrium. Suppose $a_{i}=b_{i}^{y}y+b_{i}^{x}x_{i}$ for all $i\in N$. Substituting these linear strategies into (2) yields
$b^{y}_{i}y+b^{x}_{i}x_{i}$ $=(1-d_{i}^{out})\left((1-\gamma)y+\gamma x_{i}\right)+\sum_{j\neq i}g_{ij}\left(b_{j}^{y}y+b_{j}^{x}((1-\gamma)y+\gamma x_{i})\right)$
$=\left((1-d_{i}^{out})(1-\gamma)+\sum_{j\neq i}g_{ij}b^{y}_{j}+(1-\gamma)\sum_{j\neq i}g_{ij}b^{x}_{j}\right)y+\gamma\left(1-d_{i}^{out}+\sum_{j\neq i}g_{ij}b^{x}_{j}\right)x_{i},$
where the first equality follows from $\mathbb{E}_{i}[\theta]=(1-\gamma)y+\gamma x_{i}$ and $\mathbb{E}_{i}[a_{j}]=b_{j}^{y}y+b_{j}^{x}\left((1-\gamma)y+\gamma x_{i}\right)$. By matching the coefficients, the equilibrium slopes must satisfy
$b^{y}_{i}-\sum_{j\neq i}g_{ij}b^{y}_{j}-(1-\gamma)\sum_{j\neq i}g_{ij}b^{x}_{j}$ $=(1-\gamma)(1-d_{i}^{out});$
$b^{x}_{i}-\gamma\sum_{j\neq i}g_{ij}b^{x}_{j}$ $=\gamma(1-d_{i}^{out}).$
Summing these two equations yields
$b^{y}_{i}+b^{x}_{i}-\sum_{j\neq i}g_{ij}\left(b^{y}_{j}+b^{x}_{j}\right)=1-d_{i}^{out}.$
In matrix notation,
$(I-G)(\mathbf{b}^{y}+\mathbf{b}^{x})=(I-G)\mathbf{1},$
where $\mathbf{b}^{y}=(b^{y}_{1},\ldots,b^{y}_{n})^{\prime}$ and $\mathbf{b}^{x}=(b^{x}_{1},\ldots,b^{x}_{n})^{\prime}$ are $n$-dimensional column vectors. Because the inverse matrix $(I-G)^{-1}$ exists by assumption, $\mathbf{b}^{y}+\mathbf{b}^{x}=(I-G)^{-1}(I-G)\mathbf{1}=\mathbf{1}$. Thus, any linear equilibrium must satisfy $b_{i}^{y}+b_{i}^{x}=1$ for all $i\in N$, as desired. ∎
6.2 Proposition 2
By Proposition 1, $X_{i}=(1-(1-\gamma)c_{i}(\gamma,G))\varepsilon_{i}+(1-\gamma)c_{i}(\gamma,G)\varepsilon_{y}$ does not depend on $\theta$, and the ex-ante expected payoff is well defined. I simplify $\mathbb{E}[X^{2}_{i}]$, $\mathbb{E}[X_{i}X_{j}]$, and $\sum_{j\neq i}g_{ij}\mathbb{E}[X_{i}X_{j}]$ as follows:
$\mathbb{E}[X^{2}_{i}]$ $=\mathbb{E}[((1-b_{i})\varepsilon_{i}+b_{i}\varepsilon_{y})^{2}]$
$=(1-b_{i})^{2}\tau_{x}^{-1}+b_{i}^{2}\tau_{y}^{-1}$
$=\tau_{x}^{-1}(1-(1-\gamma)c_{i}(\gamma,G))^{2}+\tau_{y}^{-1}\left((1-\gamma)c_{i}(\gamma,G)\right)^{2}$
$=(\tau_{x}^{-1}+\tau_{y}^{-1})(1-\gamma)^{2}c^{2}_{i}(\gamma,G)-2\tau_{x}^{-1}(1-\gamma)c_{i}(\gamma,G)+\tau_{x}^{-1}$
$=\tau_{x}^{-1}(1-\gamma)c^{2}_{i}(\gamma,G)-2\tau_{x}^{-1}(1-\gamma)c_{i}(\gamma,G)+\tau_{x}^{-1}$
$=\tau_{x}^{-1}(1-\gamma)c_{i}(\gamma,G)(c_{i}(\gamma,G)-2)+\tau_{x}^{-1};$
$\mathbb{E}[X_{i}X_{j}]$ $=\mathbb{E}[((1-b_{i})\varepsilon_{i}+b_{i}\varepsilon_{y})((1-b_{j})\varepsilon_{j}+b_{j}\varepsilon_{y})]$
$=b_{i}b_{j}\tau_{y}^{-1}$
$=\tau_{y}^{-1}(1-\gamma)^{2}c_{i}(\gamma,G)c_{j}(\gamma,G);$
$\sum_{j\neq i}g_{ij}\mathbb{E}[X_{i}X_{j}]$ $=\tau_{y}^{-1}(1-\gamma)^{2}c_{i}(\gamma,G)\sum_{j\neq i}g_{ij}c_{j}(\gamma,G)$
$=\tau_{y}^{-1}(1-\gamma)^{2}c_{i}(\gamma,G)\gamma^{-1}(c_{i}(\gamma,G)-1)$
$=\tau_{x}^{-1}(1-\gamma)c_{i}(\gamma,G)(c_{i}(\gamma,G)-1).$
The second last equality follows from the identity $\mathbf{c}(\gamma,G)=\mathbf{1}+\gamma G\mathbf{c}(\gamma,G)$.
Using these expressions, agent $i$’s equilibrium payoff simplifies to
$U_{i}(\tau_{x},\tau_{y},G)$ $=-\left(1-d_{i}^{out}\right)\mathbb{E}[X^{2}_{i}]-\sum_{j\neq i}g_{ij}\mathbb{E}[(X_{i}-X_{j})^{2}]$
$=-\mathbb{E}[X^{2}_{i}]-\sum_{j\neq i}g_{ij}\mathbb{E}[X_{j}^{2}]+2\sum_{j\neq i}g_{ij}\mathbb{E}[X_{i}X_{j}]$
$=\tau_{x}^{-1}(1-\gamma)\left(-c_{i}(\gamma,G)(c_{i}(\gamma,G)-2)+2c_{i}(\gamma,G)(c_{i}(\gamma,G)-1)\right)-\sum_{j\neq i}g_{ij}\mathbb{E}[X_{j}^{2}]-\tau_{x}^{-1}$
$=\tau_{x}^{-1}(1-\gamma)\left(c_{i}^{2}(\gamma,G)-\sum_{j\neq i}g_{ij}c_{j}(\gamma,G)\left(c_{j}(\gamma,G)-2\right)\right)-\tau_{x}^{-1}(1+d_{i}^{out}).$
Because $U_{i}(\tau_{x},0,G)=-\tau_{x}^{-1}(1+d_{i}^{out})$, it follows that $\Delta U_{i}(\tau_{x},\tau_{y},G)<0$ if and only if $c_{i}^{2}(\gamma,G)-\sum_{j\neq i}g_{ij}c_{j}(\gamma,G)\left(c_{j}(\gamma,G)-2\right)<0$, as desired.∎
6.3 Claims in Example 1
Claim.
$\Delta U_{A}(\tau_{x},\tau_{y},G)<0$ if and only if $f(\alpha,\beta,\gamma)<0$.
The centralities of Ann and Bob are derived using the self-referentiality: $c_{A}(\gamma,G)=1+\gamma\alpha c_{B}(\gamma,G)$, $c_{B}(\gamma,G)=1+\gamma\beta c_{C}(\gamma,G)$, and $c_{C}(\gamma,G)=1+\gamma\beta c_{B}(\gamma,G)$. Solving these simultaneous equations yields $c_{A}(\gamma,G)=(1+\gamma(\alpha-\beta))/(1-\gamma\beta)$ and $c_{B}(\gamma,G)=c_{C}(\gamma,G)=1/(1-\gamma\beta)$. For Ann, plugging these into the terms inside the large parentheses of (11) yields
$c_{A}^{2}(\gamma,G)-\alpha c_{B}(\gamma,G)\left(c_{B}(\gamma,G)-2\right)$ $=(1-\gamma\beta)^{-2}\left((1+\gamma(\alpha-\beta))^{2}-\alpha(1-2(1-\gamma\beta))\right)$
$=(1-\gamma\beta)^{-2}\left((1+\gamma(\alpha-\beta))^{2}-\alpha(2\gamma\beta-1)\right)$
$=(1-\gamma\beta)^{-2}f(\alpha,\beta,\gamma).$
Hence, (12) holds for Ann if and only if $f(\alpha,\beta,\gamma)<0$. ∎
Claim.
$\mathcal{G}(\gamma)$ is non-empty if and only if $\gamma>(1+\sqrt{5})/4$.
It suffices to show that there exists a triplet $(\alpha,\beta,\gamma)$ such that $f(\alpha,\beta,\gamma)<0$ if and only if $\gamma>(1+\sqrt{5})/4$. Define a quadratic function $h(\alpha,\gamma)\equiv\inf_{\beta\in[0,1)}f(\alpha,\beta,\gamma)$. Simple algebra shows
$h(\alpha,\gamma)$ $=f(\alpha,1,\gamma)$
$=(1+\gamma(\alpha-1))^{2}-\alpha(2\gamma-1)$
$=\gamma^{2}\alpha^{2}+2\gamma(1-\gamma)\alpha+(1-\gamma)^{2}-2\gamma\alpha+\alpha$
$=\gamma^{2}\alpha^{2}+(1-2\gamma^{2})\alpha+(1-\gamma)^{2}.$
Because $h(0,\gamma)>0$ and $h(1,\gamma)>0$ for any $\gamma\in(0,1)$, the quadratic function $h(\alpha,\gamma)$ can be strictly negative for some $\alpha\in[0,1)$ if and only if its minimum is negative. This requires two conditions to hold jointly: (i) the vertex must lie in the interval $(0,1)$, i.e., $\gamma>1/\sqrt{2}$, and (ii) the value at the vertex must be negative, i.e., $(2\gamma-1)(-4\gamma^{2}+2\gamma+1)<0$. Since $\gamma>1/\sqrt{2}$ implies that $02\gamma-1>$ 0, the two conditions are jointly equivalent to the single inequality $04\gamma^{2}-2\gamma-1>$ 0. This final inequality holds if and only if $\gamma>(1+\sqrt{5})/4$. ∎
Claim.
$\mathcal{G}(\gamma)\subseteq\mathcal{G}(\gamma^{\prime})$ for $\gamma<\gamma^{\prime}$.
This claim trivially holds if $\mathcal{G}(\gamma)$ is empty. Suppose $\mathcal{G}(\gamma)$ is nonempty. Then, $\alpha<\beta$ must hold; otherwise, $f(\alpha,\beta,\gamma)\geq 1-\alpha(2\gamma\beta-1)>0$. Differentiating $f(\alpha,\beta,\gamma)$ with respect to $\gamma$ gives
$\frac{\partial f(\alpha,\beta,\gamma)}{\partial\gamma}=2(\alpha-\beta)+2\gamma(\alpha-\beta)^{2}-2\alpha\beta=-2(\beta-\alpha)(1-\gamma(\beta-\alpha))-2\alpha\beta<0.$
Thus, $f(\alpha,\beta,\gamma)$ is strictly decreasing in $\gamma$ whenever $\alpha<\beta$. ∎
6.4 Derivation of the unique linear equilibrium under $\mathcal{I}^{\prime}$
Let Bob’s and Carol’s linear strategies be $a^{\prime}_{B}=b_{B}^{x}x_{B}+b_{B}^{y}y$ and $a^{\prime}_{C}=b_{C}^{x}x_{C}+b_{C}^{z}z$, respectively. Then, Bob’s first-order condition is
$a^{\prime}_{B}$ $=(1-\beta)(\gamma x_{B}+(1-\gamma)y)+\beta(b_{C}^{x}+b_{C}^{z})(\gamma x_{B}+(1-\gamma)y)$
$=\gamma\left(1-\beta+\beta(b_{C}^{x}+b_{C}^{z})\right)x_{B}+(1-\gamma)\left(1-\beta+\beta(b_{C}^{x}+b_{C}^{z})\right)y.$
By matching the coefficients, Bob’s equilibrium slopes satisfy
$b^{x}_{B}$ $=\gamma(1-\beta+\beta(b_{C}^{x}+b_{C}^{z}));$
$b^{y}_{B}$ $=(1-\gamma)(1-\beta+\beta(b_{C}^{x}+b_{C}^{z})).$
A similar logic applies to Carol’s equilibrium strategy, and her equilibrium slopes satisfy
$b^{x}_{C}$ $=\gamma\left(1-\beta+\beta(b_{B}^{x}+b_{B}^{y})\right);$
$b^{z}_{C}$ $=(1-\gamma)\left(1-\beta+\beta(b_{B}^{x}+b_{B}^{y})\right).$
Solving these simultaneous equations yields $b_{B}^{y}=b_{C}^{z}=1-\gamma$ and $b_{B}^{x}=b_{C}^{x}=\gamma$. Substituting the obtained Bob’s equilibrium strategy into Ann’s first-order condition gives
$a^{\prime}_{A}$ $=(1-\alpha)(\gamma x_{A}+(1-\gamma)y)+\alpha\left((1-\gamma)y+\gamma(\gamma x_{A}+(1-\gamma)y)\right)$
$=(1-\gamma)(1+\gamma\alpha)y+(\gamma(1-\alpha)+\gamma^{2}\alpha)x_{A}.$
This concludes the proof. ∎
6.5 Derivation of inequality (17)
Let $U^{\prime}_{A}(\tau_{x},\tau_{y},G)$ denote Ann’s equilibrium payoff under $\mathcal{I}^{\prime}$. By the proof in Section 6.4, Ann’s and Bob’s equilibrium strategies under $\mathcal{I}^{\prime}$ are given by $a^{\prime}_{A}=(1-\gamma)(1+\gamma\alpha)y+\gamma(1-\alpha+\gamma\alpha)x_{A}$ and $a^{\prime}_{B}=(1-\gamma)y+\gamma x_{B}$, respectively. Notice that $a^{\prime}_{A}$ and $a^{\prime}_{B}$ coincide with the equilibrium strategies under $\mathcal{I}$ for agents with centralities of $1+\gamma\alpha$ and $1$, respectively. Thus, plugging $1+\gamma\alpha$ as Ann’s centrality and $1$ as Bob’s centrality into (11) yields
$U^{\prime}_{A}(\tau_{x},\tau_{y},G)$ $=\tau_{x}^{-1}(1-\gamma)((1+\gamma\alpha)^{2}+\alpha)-\tau_{x}^{-1}(1+\alpha).$
The difference in Ann’s equilibrium payoff between $\mathcal{I}^{\prime}$ and $\mathcal{I}$ is then computed as
$U^{\prime}_{A}(\tau_{x},\tau_{y},G)-U_{A}(\tau_{x},\tau_{y},G)$ $=\tau_{x}^{-1}(1-\gamma)\left((1+\gamma\alpha)^{2}+\alpha-(1-\gamma\beta)^{-2}f(\alpha,\beta,\gamma)\right).$
Thus, $U_{A}(\tau_{x},\tau_{y},G)<U^{\prime}_{A}(\tau_{x},\tau_{y},G)$ if and only if $f(\alpha,\beta,\gamma)<(1-\gamma\beta)^{2}\left((1+\gamma\alpha)^{2}+\alpha\right)$. ∎
6.6 Proposition 3
By the proof of Proposition 2, for each $i\in N$,
$\Delta U_{i}(\tau_{x},\tau_{y},G)$ $=\tau_{x}^{-1}(1-\gamma)\left(c_{i}^{2}(\gamma,G)-\sum_{j\neq i}g_{ij}c_{j}(\gamma,G)\left(c_{j}(\gamma,G)-2\right)\right).$
Summing these equations over $i$ yields
$\Delta W(\tau_{x},\tau_{y},G)$ $=\sum_{i\in N}\Delta U_{i}(\tau_{x},\tau_{y},G)$
$=\tau_{x}^{-1}(1-\gamma)\left(\sum_{i\in N}c_{i}^{2}(\gamma,G)-\sum_{i\in N}\sum_{j\neq i}g_{ij}c_{j}(\gamma,G)\left(c_{j}(\gamma,G)-2\right)\right)$
$=\tau_{x}^{-1}(1-\gamma)\left(\sum_{i\in N}c_{i}^{2}(\gamma,G)-\sum_{i\in N}d_{i}^{in}c_{i}(\gamma,G)\left(c_{i}(\gamma,G)-2\right)\right)$
$=\tau_{x}^{-1}(1-\gamma)\sum_{i\in N}c_{i}(\gamma,G)\left((1-d_{i}^{in})c_{i}(\gamma,G)+2d_{i}^{in}\right).$
Hence, $\Delta W(\tau_{x},\tau_{y},G)<0$ if and only if $\sum_{i\in N}c_{i}(\gamma,G)\left((1-d_{i}^{in})c_{i}(\gamma,G)+2d_{i}^{in}\right)<0$. ∎
6.7 Claims in Example 2
Claim.
Suppose $G$ is a core-periphery interaction network. Then, $\Delta W(\tau_{x},\tau_{y},G)<0$ if and only if $f(\alpha,\beta,\gamma)<-l(1-\beta(2\gamma\beta-1))/m$.
Suppose $G$ is a core-periphery interaction network. Each agent $i$’s centrality is given by $c_{i}(\gamma,G)=1/(1-\gamma\beta)$ if $i$ is core and $c_{i}(\gamma,G)=(1+\gamma(\alpha-\beta))/(1-\gamma\beta)$ if $i$ is peripheral. The sum of all core agents’ in-degrees is $l\beta+m\alpha$, and the sum of all peripheral agents’ in-degrees is 00. Plugging these into the left-hand side of (18) yields
$\mathrm{(LHS)}$ $=(1-\gamma\beta)^{-2}\left(l-(l\beta+m\alpha)+2(1-\gamma\beta)(l\beta+m\alpha)+m(1+\gamma(\alpha-\beta))^{2}\right)$
$=(1-\gamma\beta)^{-2}\left(l+(l\beta+m\alpha)(1-2\gamma\beta)+mf(\alpha,\beta,\gamma)+m\alpha(2\gamma\beta-1)\right)$
$=(1-\gamma\beta)^{-2}\left(mf(\alpha,\beta,\gamma)+l(1-\beta(2\gamma\beta-1)\right)).$
Hence, (18) holds if and only if $f(\alpha,\beta,\gamma)<-l(1-\beta(2\gamma\beta-1))/m$. ∎
Claim.
$\mathcal{H}(\gamma,q)\subseteq\mathcal{H}(\gamma,p)$ for $p\leq q$.
This is immediate from the fact that $01-\beta(2\gamma\beta-1)>$ 0. ∎
Claim.
For each $\gamma\in(0,1)$, $\mathcal{H}(\gamma,l/m)\rightarrow\mathcal{G}(\gamma)$ as $l/m\rightarrow 0$.
Fixing $\gamma\in(0,1)$, the right-hand side of (19) converges to 0 as $l/m\rightarrow 0$. The obtained inequality in the limit coincides with inequality (13). ∎
6.8 Proposition 4
To show this proposition, it suffices to focus on the three-player interaction network in Example 1 and show that for any $\gamma>1/\sqrt{2}$, there exist interaction parameters $(\alpha,\beta)\in[0,1)^{2}$ such that
$\frac{\partial f}{\partial\alpha}<0.$
Specifically, the partial derivative is computed as
$\frac{\partial f}{\partial\alpha}=2\gamma+2\gamma^{2}(\alpha-\beta)-2\gamma\beta+1=2\gamma^{2}\alpha-2\gamma(\gamma+1)\beta+2\gamma+1.$
The infimum of $\partial f/\partial\alpha$ with respect to $(\alpha,\beta)$ is then
$\inf_{(\alpha,\beta)\in[0,1)^{2}}\frac{\partial f}{\partial\alpha}=-2\gamma^{2}+1.$
Because $-2\gamma^{2}+1<0$ if and only if $\gamma>1/\sqrt{2}$, the result follows. ∎
6.9 Derivation of the unique linear equilibrium under $\mathcal{I}^{\dagger}$
Following the proof in Section 6.1, it is easy to show that the sum of equilibrium slopes must be one for each agent in any linear equilibrium under $\mathcal{I}^{\dagger}$. This observation immediately implies that agent $i\neq 1$’s unique linear equilibrium strategy must be $a^{\dagger}_{i}=x_{i}$. Plugging these into agent $1$’s first-order condition yields
$a^{\dagger}_{1}=(1-d_{1}^{out})\mathbb{E}_{1}[\theta]+\sum_{j\neq 1}g_{1j}\mathbb{E}_{1}[x_{j}]=\mathbb{E}_{1}[\theta]=\gamma x_{1}+(1-\gamma)y.$
This concludes the proof. ∎
6.10 Derivation of $U^{\dagger}_{i}(\tau_{x},\tau_{y},G)$ and $W^{\dagger}(\tau_{x},\tau_{y},G)$
First, I derive agent $1$’s equilibrium payoff under $\mathcal{I}^{\dagger}$:
$U^{\dagger}_{1}(\tau_{x},\tau_{y},G)$ $=-\left(1-d_{1}^{out}\right)\mathbb{E}[(a^{\dagger}_{1}-\theta)^{2}]-\sum_{j\neq 1}g_{1j}\mathbb{E}[(a^{\dagger}_{1}-a^{\dagger}_{j})^{2}]$
$=-\mathbb{E}[(a^{\dagger}_{1}-\theta)^{2}]-\sum_{j\neq 1}g_{1j}\mathbb{E}[(a^{\dagger}_{j}-\theta)^{2}]+2\sum_{j\neq 1}g_{1j}\mathbb{E}[(a^{\dagger}_{1}-\theta)(a^{\dagger}_{j}-\theta)]$
$=-\mathbb{E}[(a^{\dagger}_{1}-\theta)^{2}]-\sum_{j\neq 1}g_{1j}\mathbb{E}[(a^{\dagger}_{j}-\theta)^{2}]$
$=-(1-\gamma)^{2}\tau_{y}^{-1}-\gamma^{2}\tau_{x}^{-1}-d_{1}^{out}\tau_{x}^{-1}$
$=-\gamma\tau_{x}^{-1}-d_{1}^{out}\tau_{x}^{-1}$
$=\tau_{x}^{-1}(1-\gamma)-\tau_{x}^{-1}(1+d_{1}^{out}).$
Similarly, agent $i\neq 1$’s equilibrium payoff under $\mathcal{I}^{\dagger}$ is derived as
$U^{\dagger}_{i}(\tau_{x},\tau_{y},G)$ $=-\tau_{x}^{-1}-\sum_{j\neq i,1}g_{ij}\tau_{x}^{-1}-g_{i1}((1-\gamma)^{2}\tau_{y}^{-1}+\gamma^{2}\tau_{x}^{-1})$
$=\tau_{x}^{-1}(1-\gamma)g_{i1}-\tau_{x}^{-1}(1+d_{i}^{out}).$
Thus, $U^{\dagger}_{1}(\tau_{x},\tau_{y},G)=\tau_{x}^{-1}(1-\gamma)+U_{1}(\tau_{x},0,G)$ and $U^{\dagger}_{i}(\tau_{x},\tau_{y},G)=\tau_{x}^{-1}(1-\gamma)g_{i1}+U_{i}(\tau_{x},0,G)$ for each $i\neq 1$. Summing over all agents then yields $W^{\dagger}(\tau_{x},\tau_{y},G)=\tau_{x}^{-1}(1-\gamma)(1+d^{in}_{1})+W(\tau_{x},0,G)$, which concludes the proof. ∎
6.11 Proposition 5
By the proof of Proposition 2 and Proposition 3,
$U^{\dagger}_{1}(\tau_{x},\tau_{y},G)-U_{1}(\tau_{x},\tau_{y},G)$ $=\tau_{x}^{-1}(1-\gamma)+U_{1}(\tau_{x},0,G)-U_{1}(\tau_{x},\tau_{y},G)$
$=\tau_{x}^{-1}(1-\gamma)-\Delta U_{1}(\tau_{x},\tau_{y},G)$
$=\tau_{x}^{-1}(1-\gamma)\left(1-c_{1}^{2}(\gamma,G)+\sum_{j\neq 1}g_{1j}c_{j}(\gamma,G)\left(c_{j}(\gamma,G)-2\right)\right);$
$W(\tau_{x},\tau_{y},G)-W^{\dagger}(\tau_{x},\tau_{y},G)$ $=W(\tau_{x},\tau_{y},G)-\tau_{x}^{-1}(1-\gamma)(1+d^{in}_{1})-W(\tau_{x},0,G)$
$=\Delta W(\tau_{x},\tau_{y},G)-\tau_{x}^{-1}(1-\gamma)(1+d^{in}_{1})$
$=\tau_{x}^{-1}(1-\gamma)\left(\sum_{i\in N}c_{i}(\gamma,G)\left((1-d_{i}^{in})c_{i}(\gamma,G)+2d_{i}^{in}\right)-(1+d^{in}_{1})\right).$
Thus, $U_{1}(\tau_{x},\tau_{y},G)<U^{\dagger}_{1}(\tau_{x},\tau_{y},G)$ and $W^{\dagger}(\tau_{x},\tau_{y},G)<W(\tau_{x},\tau_{y},G)$ if and only if $c_{1}^{2}(\gamma,G)-\sum_{j\neq 1}g_{1j}c_{j}(\gamma,G)\left(c_{j}(\gamma,G)-2\right)<1$ and $1+d^{in}_{1}<\sum_{i\in N}c_{i}(\gamma,G)\left((1-d_{i}^{in})c_{i}(\gamma,G)+2d_{i}^{in}\right)$. ∎
6.12 Derivation of compound inequality (25)
Suppose $G$ is a core-periphery interaction network. By substituting the simplified expressions for the left-hand side of (23) (derived in Section 6.3) and the right-hand side of (24) (derived in Section 6.7), the two conditions can be combined into the single compound inequality:
$(1-\gamma\beta)^{-2}f(\alpha,\beta,\gamma)<1<(1-\gamma\beta)^{-2}\left(mf(\alpha,\beta,\gamma)+l(1-\beta(2\gamma\beta-1))\right).$
Multiplying all parts of this inequality by $(1-\gamma\beta)^{2}$ yields the compound inequality (25). ∎